OpenStudy (kayla2001):

y=2x-1 and y=x+3 a.) 4,7 b.) 7,4 c.) -7,-4 d.) no solutions

OpenStudy (kayla2001):

@ igreen I need help please

OpenStudy (clongoria2):

Just plugin those points into both equations and see if one point equals for both equation.

OpenStudy (kayla2001):

I need somebody please

syed98 (syedmohammed98):

You can use the elimination method :)

OpenStudy (clongoria2):

Ok so ill break it down for you

OpenStudy (kayla2001):

I thought it was b

OpenStudy (clongoria2):

4,7 4 = x 7 = y. The equations are y = 2x -1 and y = x + 3 So plugin 4,7 to both equations 7 = 2 x 4 - 1 and 7 = 4 + 3 8 - 1 = 7 and 7 = 4 + 3 = 7

OpenStudy (kayla2001):

so it id d

OpenStudy (clongoria2):

@iGreen. @AnswerMyQuestions Maybe you guys will be able to explain it to her better....

OpenStudy (clongoria2):

There is a solution. Ill give you that much....

OpenStudy (kayla2001):

ooooooooo its a

OpenStudy (clongoria2):

Do you understand why?

OpenStudy (kayla2001):

yes maim or sir

OpenStudy (anonymous):

y = 2x - 1 y = x + 3 We can substitute 2x - 1 for 'y' in the 2nd equation: 2x - 1 = x + 3 Add 1 to both sides, what's 3 + 1?

OpenStudy (anonymous):

Btw, you're not supposed to give out the answer. @Clongoria2

OpenStudy (clongoria2):

@iGreen. I know, but she wasn't understanding it well. Sorry